# GPU-based compressible lattice Boltzmann simulations on non-uniform grids using standard C++ parallelism: From best practices to aerodynamics, aeroacoustics and supersonic flow simulations

### v2

@@ 9 General conclusions and perspectives
The physical validation of our framework relies on an extensive set of benchmarks, covering a wide range of flow regimes from low-speed aerodynamics and aeroacoustic to supersonic polyatomic flows. These benchmarks include simulations of lid-driven cavity flows, Aeolian noise generated by the flow past a circular cylinder, external aerodynamics past the 30P30N airfoil at realistic conditions, inviscid Riemann problems, as well as, the viscous flow past a NACA airfoil in both transonic and supersonic regimes. These validations show that:
  • Weakly compressible LBMs based on modern collision models, with a conservative CC grid refinement strategy and simple boundary conditions, can accurately and robustly simulate fluid flows for Reynolds numbers ranging from 100 to over $10^{6}$, and for moderate local Mach numbers up to 0.5-0.6.
  • Simple extensions of the grid refinement strategy to high-order LBMs achieve accuracy and robustness comparable to the original approach. These findings hold across transonic and supersonic regimes, in both inviscid and viscous conditions. Notably, achieving such accuracy in inviscid and under-resolved scenarios represents a significant milestone in the field of supersonic flow simulations with pure LBMs.
  • Our refinement strategy is general enough to seamlessly handle refinement interfaces intersecting inlet, outlet, and no-slip boundary conditions. This allows for refinement patches with complex shapes that keep the memory footprint, and the corresponding simulation time to a minimum.
Overall, the validation results demonstrate that CC-based grid refinement strategies provide an excellent balance of accuracy, robustness, and efficiency for both weakly and fully compressible LBMs. Moreover, their inherent conservation of mass, momentum, and energy offers a clear advantage over the more widely used vertex-centered (VC) approaches. Their versatility also allows for seamless extension to adaptive mesh refinement (AMR), with preliminary validations conducted on convected vortices, double shear layers, and supersonic flows past a NACA airfoil using D2Q21-LBMs. These observations undoubtedly reposition Rohde’s refinement approach, and similar strategies, at the forefront of discussions on the design of industry-oriented LBMs.
To implement LBMs that efficiently run on non-uniform grids, we derived performance models that serve as reference for optimal efficiency on a given hardware. For the first time, these models were derived by incorporating key components of the grid refinement strategy, namely, the neighbor-list for streaming in tree-like grids, and the asynchronous time-stepping. Where possible, theoretical performance models were compared to practical measurements from lid-driven cavity, Aeolian noise, and Sod shock tube simulations. Our main findings are:
  • The list of neighbors requires additional memory accesses, which cuts performance approximately in half for both weakly and fully compressible LBMs. Moreover, if not properly anticipated, the combination of GPU memory underuse and asynchronous time-stepping can significantly degrade performance. However, the latter can also greatly reduce the simulation time if the user adjust carefully the number of cells in the levels with the highest update frequency, while keeping a fixed geometry resolution.
  • The C++ implementation of weakly compressible LBMs, with tree-based grid refinement, exhibits raw performance comparable to CUDA implementations in state-of-the-art solvers like waLBerla. Fully compressible LBMs, being more compute-intensive, show lower raw performance than memory-bound solvers but achieve greater GPU speedups.
@@
Overall, the performance analysis demonstrates that ISO languages, like C++, now compete with low-level, hardware-specific languages in performance while offering superior modularity and ease of implementation. As such, they are viable alternatives for rapid prototyping, as well as more advanced design of accurate, robust, and efficient LBMs on non-uniform grids.
In conclusion, this work establishes a new baseline for developing GPU-accelerated, tree-based LBMs for weakly and fully compressible flows. It demonstrates that most academic applications do not require complex LBMs, boundary conditions, or grid refinement strategies to achieve state-of-the-art results. However, this framework requires further improvements to reach the standards of industrial solvers. The following modifications are currently being investigated and will be part of future publications:
  • Memory footprint: Grids composed of billions of fluid cells are becoming more common in the industry, but the most affordable GPU cards can hardly store the data of more than half a billion cells using single precision. As such, the number of variables stored in memory should be kept to a minimum to run large simulations on limited computational resources. Here, the “two-population strategy” (one for reads and another for writes) was used for rapid prototyping of advanced LBMs, but it doubles the memory storage requirements compared to in-place implementations of the streaming step. Among these approaches, out attention will focus on the AA pattern [206] which can halve the memory footprint without drastically increasing code complexity.
  • Adaptive mesh refinement: While this work provides preliminary proofs of the accuracy and robustness of our CPU-based AMR strategy, efficient mesh generation and adaptation on the GPU are necessary to perform realistic simulations on affordable hardware. Developing GPU-based dynamic mesh refinement strategies, driven by flow-dependent or kinetic criteria [173], would allow for real-time adaptation and significantly enhance the solver’s performance for multi-scale flow problems.
@@
  • Boundary conditions: Staircased boundary conditions, like the present bounce-back approach, are known to suffer from accuracy issues in under-resolved conditions [85]. Moving to curved boundary conditions that account for the distance to the wall (e.g., Bouzidi et al.[207] or Filippova and Hänel [15]) is a necessary step to efficiently run simulations in under-resolved conditions while ensuring a good level of accuracy close to the wall.
@@
  • Wall modeling: The combination of curved boundary conditions with wall models, such as algebraic wall functions or more advanced approaches, is crucial for accurately capturing near-wall effects in high-Reynolds-number flows [208]. Integrating such models is particularly important for achieving the required accuracy without fully resolving boundary layer development [209]. This ensures computational efficiency while making the solver more suitable for industrial applications.
@@
  • Turbulence modeling: While the numerical dissipation introduced by the collision model can stabilize simulations at high Reynolds numbers, explicit turbulence models are essential for accurately capturing dissipation mechanisms that cannot be fully resolved by coarse grids. Implementing subgrid-scale models (Smagorinsky, Vreman, or WALE models) would enable more accurate simulations at coarse grid resolutions, while allowing for a better control over the amount of turbulent dissipation introduced in the simulation domain [210, 152].
These aspects are progressively being incorporated in a 3D version of the present framework. Corresponding results will be presented in future work dedicated to multi-GPU, industry-oriented 3D-LBM with adaptive mesh refinement.
@@ Acknowledgements
[…]
@@ A.1 General knowledge
where $f^{*}_{i}$ are post collision populations. Eventually, we recall in Tables 4 and 5 the properties of the velocity discretizations used in this work.
@@ A.2 Advanced models for accurate and robust simulations of low and high speed flows
Here, $\Pi_{\alpha\beta}=-\rho c_{s}^{4}\tau\left(\partial_{\alpha}u_{\beta}+\partial_{\beta}u_{\alpha}-\textstyle{\frac{2}{D}}\partial_{\gamma}u_{\gamma}\right)\approx\textstyle{\sum_{i}}\mathcal{H}^{(2)}_{\alpha\beta}\left(f_{i}-f_{i}^{eq}\right)$. $\Pi_{\alpha\beta}$ can be computed using either LB data, finite differences, or a combination of both for enhanced stability [76]. Here, $\Pi_{\alpha\beta}$ is computed using only LB data. If additional robustness is required for purely aerodynamics simulations, equilibrating the trace of second-order (Hermite) moments is usually a good choice. This approach is widely used in moment-based collision models by setting the bulk viscosity relaxation time to unity [143, 144, 145, 146, 147, 148, 149]. To ensure consistency and avoid case-by-case deliberations in Section 5, we systematically apply this condition in aerodynamic simulations, while excluding it for aeroacoustic simulations.
@@ A.4 Known defects of grid refinement strategies for computational aeroacoustics
  • CC approach with uniform explosion and a buffer layer: Rohde et al. [24] proposed the CC approach with uniform explosion, which has been implemented in the open-source solver waLBerla [100]. This method is referred to as ‘cc (uni)’.
  • CC with linear explosion: Extending cc (uni)’, Chen et al. [25] introduced the linear explosion that accounts for local gradients of populations. This approach has been implemented in the commercial software PowerFLOW [33] and is referred to as ‘cc (lin)’.
@@ B.2 Lid-driven cavity simulations at higher Reynolds numbers
The fluid flow within the square lid-driven cavity becomes unsteady when the Reynolds number reaches the critical value of $8000$, a phenomenon known as Hopf bifurcation [113, 114]. Therefore, we simulate this benchmark at $\mathrm{Re}=7500$, the highest Reynolds number for which steady-state reference data is available [116]. At this higher Reynolds number, the boundary layer developing along the walls is much thinner than at $\mathrm{Re}=100$, hence, refinement patches should be added accordingly. Consequently, the vert mesh configuration is no longer suitable, and another mesh composed of three nested refinement layers (3 lay) is adopted instead. Figures 30 and 31 display the meshes used for this configuration and the velocity profiles along the cavity centerlines, respectively. These profiles, plotted for the D2Q9, D2Q13, and D2Q21 lattices, show excellent agreement with the reference data.
@@ B.3 Aeolian noise: Impact of Reynolds and Mach numbers
In this study, additional simulations were performed to investigate the influence of Reynolds and Mach numbers on the noise generated by the flow past a circular cylinder. Two parametric studies were conducted, and they are dedicated to: (1) examining the effect of varying the Reynolds number while keeping the Mach number constant at $\mathrm{Ma}=0.2$, and (2) analyzing the effect of varying the Mach number while maintaining a fixed Reynolds number of $\mathrm{Re}=150$. The second study also provides the opportunity to evaluate the accuracy and robustness of the refinement strategy near the stability limit of the D2Q9-RR-LBM. Notably, the case with $\mathrm{Ma}=0.5$ results in a local Mach number of approximately 0.72, approaching the theoretical stability threshold for weakly compressible LBMs, $1-\sqrt{3}\approx 0.73$ [212, 243, 244, 75, 7]. Corresponding pressure fluctuation fields are shown in Figures 32 and 33.
@@ B.6 Transonic NACA flow: Preliminary simulations without local refinement at the leading edge
On top of the performance analyses conducted with the D2Q9 lattice for several collision models (BGK, TRT and RR), we provide additional data corresponding to two high-order lattices, namely, the D2Q13 and D2Q21. First, we checked the implementation of these high-order LBMs on a matrix-like data structure, with SoA memory layout, and “two-population” strategy. This implementation is based on the STLBM framework [11] and served as reference before including the influence of the list of neighbors required for tree-like data structures.
Figures 41 compares results obtained with the three collision models, as well as single and double precision arithmetics. As for the performance on D2Q9-LBMs (Section 7.2), all configuration reach 80% of the peak performance, $P_{\mathrm{peak}}^{\mathrm{matrix}}$, which is 18.21 (9.44) and 11.59 (5.93) for single (double) precision with the D2Q13 and D2Q21 LBMs, respectively.
@@
Figure 42 presents results for the same configuration but utilizing tree-like data structures. Due to the partial optimization of the code, the high-order RR-LBMs achieve only 45% of the peak performance, $P_{\mathrm{peak}}^{\mathrm{matrix}}$, which corresponds to 12.74 (7.72) and 7.96 (4.81) GLUPS for single (double) precision with the D2Q13 and D2Q21 lattices, respectively. The fact that performance varies with the collision model, despite the LBMs being memory-bound, also highlights the incomplete optimization of the collision step. Nonetheless, this percentage remains close to highly optimized LBMs implemented in waLBerla using trees of blocks and CUDA/HIP [10].
@@
[…]
@@
The consistent findings across all finest grid resolutions ($\Delta x_{f}\in\{D/20,D/40,D/80\}$) underline the importance of tailoring the grid to the specific simulation objectives. Whether the focus is on near-field aerodynamics or far-field acoustics, adopting best practices to balance accuracy, performance, and simulation time remains crucial. The performance model developed here and the insights from the case studies offer a robust methodology for optimizing grid generation in a semi-automated manner, even for more complex scenarios requiring large domains or localized refinement patches.

### v3

@@ 9 Synthesis of findings
Our framework is validated through an extensive set of 2D benchmarks, covering a wide range of flow regimes from low-speed aerodynamics and aeroacoustic to supersonic polyatomic flows. These benchmarks include simulations of lid-driven cavity flows, Aeolian noise generated by the flow past a circular cylinder, external aerodynamics past the 30P30N airfoil at realistic conditions, inviscid Riemann problems, as well as, the viscous flow past a NACA airfoil in both transonic and supersonic regimes. These validations were designed to address two central questions:
  1. 1.
  2. 2.
The results provide affirmative answers to both questions and further reveal that:
  • The inherent conservation properties of CC-based grid refinement strategies offer a clear advantage over the more widely used vertex-centered (VC) approaches, especially in the context of aeroacoustics simulations.
  • The refinement strategy naturally handles complex geometries, including interfaces intersecting inlets, outlets, and no-slip boundaries, allowing for flexible mesh designs with minimal memory and computational overhead.
  • Its versatility also allows for seamless extension to adaptive mesh refinement (AMR), with preliminary validations conducted on convected vortices, double shear layers, and supersonic flows past a NACA airfoil using D2Q21-LBMs.
Overall, the proposed approach combines simplicity, generality, and robustness which allows for accurate LB simulations on non-uniform grids across a wide spectrum of flow conditions.
Performance study
To evaluate the computational efficiency of our framework on non-uniform grids, we focused on answering the following two questions:
  1. 1.
  2. 2.
To answer the first question, we derived analytical performance models that incorporate extra memory accesses to the neighbor-list, as well as, the asynchronous time-stepping though a level-wise weighted performance model. Where possible, theoretical performance models were compared to practical measurements from lid-driven cavity, Aeolian noise, and Sod shock tube simulations. Notably, we found out that:
  • The overhead induced by the neighbor-list cuts performance approximately in half for both weakly and fully compressible LBMs.
  • If the update frequency is similar across all refinement levels, the speedup achieved through asynchronous time-stepping is equivalent to the ratio of the total number of cells to the number of equivalent fine cells.
  • If not properly anticipated, asynchronous time-stepping can significantly degrade raw performance due to GPU memory underuse or imbalanced memory accesses between levels. Nevertheless, the resulting reduction in simulation time often compensates for the raw performance drop, yielding substantial overall speedups.
  • Our C++ implementation of weakly compressible LBMs with tree-based refinement achieves raw performance comparable to state-of-the-art CUDA-native solvers.
  • Fully compressible LBMs with numerical equilibria have lower raw performance due to their higher computational intensity. However, they achieve better GPU speedups, even matching the performance of 1000 CPU cores in the most compute-intensive scenarios.
@@
In summary, our performance analysis is able to properly isolate the contribution of each element of the grid refinement algorithm. Furthermore, it shows that ISO-standard languages like C++ can rival hardware-specific approaches in raw efficiency while offering superior ease of implementation, modularity and maintainability. This positions modern C++ as a compelling choice not only for rapid prototyping but also for more advanced design of accurate, robust, and efficient LBMs on non-uniform grids.
General conclusion
This work shows that the combination of simple LBMs, boundary conditions, and grid refinement strategies can lead to excellent results in the 2D context. Furthermore, it repositions Rohde’s refinement approach, and similar CC strategies, at the forefront of discussions on the design of industry-oriented LBMs. Thanks to modern C++, the GPU port of LBMs is rather straightforward while showing performance comparable to GPU-native solvers. In the end, this study establishes a new baseline for developing GPU-accelerated, tree-based LBMs for weakly and fully compressible flows.
10 Current and future directions
Looking ahead, we plan to enhance the current framework to meet the rigorous demands of 3D industrial solvers. To achieve this, we are investigating several key aspects:
  • Memory footprint: As grids composed of billions of fluid cells become more common in the industry, memory constraints on affordable GPU cards pose a significant challenge. These cards often struggle to store 3D data for more than half a billion cells using single precision. Thus, minimizing memory usage is crucial. While our current “two-population” strategy simplifies implementation, it also doubles memory consumption. We are exploring more memory-efficient alternatives, such as the AA pattern [207], which can potentially halve the memory footprint with minimal added complexity.
  • Multi-GPU capabilities: To scale simulations across multiple GPUs, efficient domain decomposition and communication schemes are essential, as they enable handling larger and more complex 3D cases. Although the present study focuses on single-GPU performance, recent advancements reported in our GPU port of the Palabos fluid solver show promising results for multi-GPU scalability, albeit currently limited to uniform grids [208]. Extending these strategies to support grid refinement and adaptive meshes remains to be done.
  • Adaptive mesh refinement: While initial CPU-based results have shown the potential of our AMR strategy, developing GPU-efficient implementations remains a priority. We aim to enable dynamic mesh adaptation on-the-fly using flow-dependent or kinetic criteria [173], thereby improving accuracy and efficiency for multi-scale problems.
@@ 10 Current and future directions
  • Boundary conditions: Staircased boundary conditions, such as the current bounce-back approach, are known to suffer from accuracy issues in under-resolved conditions [85]. Moving to curved boundary conditions that account for the distance to the wall (e.g., Bouzidi et al. [209] or Filippova and Hänel [15]) is a necessary step to ensure accurate simulations in under-resolved conditions while maintaining a high level of accuracy near the wall.
@@
  • Wall modeling: Combining curved boundary conditions with wall models, such as algebraic wall functions or more advanced approaches, is crucial for accurately capturing near-wall effects in high-Reynolds-number flows [210, 3]. Integrating such models is particularly important for achieving the required accuracy without fully resolving the boundary layer, ensuring computational efficiency and suitability for industrial applications [211].
@@
  • Turbulence modeling: While the numerical dissipation introduced by the collision model can stabilize simulations at high Reynolds numbers, explicit turbulence models are essential for accurately accounting for dissipation mechanisms that cannot be fully resolved by coarse grids. Implementing subgrid-scale models (Smagorinsky, Vreman, or WALE models) would enable more accurate simulations at coarse grid resolutions, while allowing for better control over the amount of turbulent dissipation introduced in the simulation domain [212, 152].
These enhancements are being progressively integrated into a 3D version of the framework. Figure 28 gathers preliminary results about the flow past a sphere that is simulated with the D3Q27-RR in the super-critical regime using AMR and the kinetic criterion. More details will be presented in future work dedicated to multi-GPU, industry-oriented 3D-LBM with AMR capabilities.
Fig 28: Illustration of AMR for the flow past a sphere with the D3Q27-$R$-LBM in the super-critical regime $(\mathrm{Re},\mathrm{Ma})=(10^{6},0.2)$. Left: Local Mach number. Right: Dynamic grid. Seven refinement levels are used to discretize the $48D\times 32D\times 32D$ simulation domain with $D$ being the sphere diameter. Equilibration of the trace of second-order moments is used for improved robustness. The finest mesh resolution is set to $\Delta x_{f}=D/200$, a threshold value required to get the flow separation past the sphere’s maximum diameter –which is characteristic of the supercritical regime.
@@ Acknowledgements
C.C. is grateful to Alexander Schukmann for insightful discussions on grid refinement strategies and for sharing an early version of Ref. [14]. Additionally, C.C. thanks Markus Holzer for providing valuable insights into the impact of grid refinement on the GPU performance of LBMs and for discussing strategies to mitigate these effects. Eventually, the authors would also like to thank the reviewers for their pertinent comments that help us improving the quality of the manuscript. This work received financial support from BNBU Research Fund $\mathrm{n^{o}UICR0700094}$-$24$ entitled “GPU-accelerated multidisciplinary methods for industrial computational fluid dynamics”.
@@ A.1 General knowledge
where $f^{*}_{i}$ are post collision populations. Eventually, we recall in Tables 5 and 6 the properties of the velocity discretizations used in this work.
@@ A.2 Advanced models for accurate and robust simulations of low and high speed flows
Here, $\Pi_{\alpha\beta}=-\rho c_{s}^{4}\tau\left(\partial_{\alpha}u_{\beta}+\partial_{\beta}u_{\alpha}-\textstyle{\frac{2}{D}}\partial_{\gamma}u_{\gamma}\right)\approx\textstyle{\sum_{i}}\mathcal{H}^{(2)}_{\alpha\beta}\left(f_{i}-f_{i}^{eq}\right)$. $\Pi_{\alpha\beta}$ can be computed using either LB data, finite differences, or a combination of both for enhanced stability [76]. Here, $\Pi_{\alpha\beta}$ is computed using only LB data. If additional robustness is required for purely aerodynamics simulations, equilibrating the trace of second-order (Hermite) moments is usually a good choice. This approach is widely used in moment-based collision models by setting the bulk viscosity relaxation time to unity [143, 144, 145, 146, 147, 148, 149]. To ensure consistency and avoid case-by-case deliberations in Section 5, we systematically apply this condition in aerodynamic simulations, while excluding it for aeroacoustic simulations. The 3D simulation of the turbulent flow past a sphere is based on the same framework, with more details available in Ref. [72].
@@ A.4 Known defects of grid refinement strategies for computational aeroacoustics
  • CC approach with uniform explosion and a buffer layer: Rohde et al. [24] proposed the CC approach with uniform explosion, which has been implemented in the open-source solver waLBerla [100]. This method is referred to as ‘$c$ (uni)’.
  • $C$ with linear explosion: Extending $c$ (uni)’, Chen et al. [25] introduced the linear explosion that accounts for local gradients of populations. This approach has been implemented in the commercial software PowerFLOW [33] and is referred to as ‘$c$ (lin)’.
@@ B.2 Lid-driven cavity simulations at higher Reynolds numbers
The fluid flow within the square lid-driven cavity becomes unsteady when the Reynolds number reaches the critical value of $8000$, a phenomenon known as Hopf bifurcation [113, 114]. Therefore, we simulate this benchmark at $\mathrm{Re}=7500$, the highest Reynolds number for which steady-state reference data is available [116]. At this higher Reynolds number, the boundary layer developing along the walls is much thinner than at $\mathrm{Re}=100$, hence, refinement patches should be added accordingly. Consequently, the vert mesh configuration is no longer suitable, and another mesh composed of three nested refinement layers (3 lay) is adopted instead. Figures 31 and 32 display the meshes used for this configuration and the velocity profiles along the cavity centerlines, respectively. These profiles, plotted for the D2Q9, D2Q13, and D2Q21 lattices, show excellent agreement with the reference data.
@@ B.3 Aeolian noise: Impact of Reynolds and Mach numbers
In this study, additional simulations were performed to investigate the influence of Reynolds and Mach numbers on the noise generated by the flow past a circular cylinder. Two parametric studies were conducted, and they are dedicated to: (1) examining the effect of varying the Reynolds number while keeping the Mach number constant at $\mathrm{Ma}=0.2$, and (2) analyzing the effect of varying the Mach number while maintaining a fixed Reynolds number of $\mathrm{Re}=150$. The second study also provides the opportunity to evaluate the accuracy and robustness of the refinement strategy near the stability limit of the D2Q9-RR-LBM. Notably, the case with $\mathrm{Ma}=0.5$ results in a local Mach number of approximately 0.72, approaching the theoretical stability threshold for weakly compressible LBMs, $1-\sqrt{3}\approx 0.73$ [214, 245, 246, 75, 7]. Corresponding pressure fluctuation fields are shown in Figures 33 and 34.
@@ B.6 Transonic NACA flow: Preliminary simulations without local refinement at the leading edge
On top of the performance analyses conducted with the D2Q9 lattice for several collision models (BGK, TRT and RR), we provide additional data corresponding to two high-order lattices, namely, the D2Q13 and D2Q21. First, we checked the implementation of these high-order LBMs on a matrix-like data structure, with SoA memory layout, and “two-population” strategy. This implementation is based on the STLBM framework [11]. Figures 42 compares results obtained with the three collision models, as well as single and double precision arithmetics. As for the performance on D2Q9-LBMs (Section 7.2), all configuration reach 80% of the peak performance, $P_{\mathrm{peak}}^{\mathrm{matrix}}$, which is 18.21 (9.44) and 11.59 (5.93) for single (double) precision with the D2Q13 and D2Q21 LBMs, respectively.
@@
To isolate the impact of the neighbor-list on performance and to compare it with our theoretical predictions ($P_{\mathrm{peak}}^{\mathrm{tree}}$), we performed additional simulations using a tree-like structure while retaining the same uniform configuration. The corresponding results are shown in Figure 43. Due to partial code optimization, the high-order RR-LBMs reach 45% of $P_{\mathrm{peak}}^{\mathrm{tree}}$. This corresponds to 12.74 (7.72) and 7.96 (4.81) GLUPS in single (double) precision for the D2Q13 and D2Q21 lattices, respectively. The variation in performance across collision models, despite the memory-bound nature of these schemes, further underlines the incomplete optimization of the collision step.
Lastly, by comparing the measured performance between matrix-based and tree-like structures, we can assess the overhead introduced by the neighbor list. In the most optimized configuration, based on the RR collision model, the performance ratio ranges from 0.38 to 0.46 (see Table 7). This indicates that employing neighbor lists reduces performance to a bit more than half of an equivalent solver operating on uniform grids. This observation, once again, aligns with results from highly optimized LBMs which rely on trees of blocks and GPU-specific languages [10].
Single Precision (GLUPS) Double Precision (GLUPS)
Lattice Matrix Neighbor $P_{\mathrm{ratio}}$ Matrix Neighbor $P_{\mathrm{ratio}}$
Q9 20.584 10.541 0.51 10.795 7.041 0.65
Q13 14.926 5.634 0.38 7.561 3.452 0.46
Q21 9.003 3.405 0.38 4.716 2.140 0.45
Table 7: Measured performance ratios $P_{\mathrm{ratio}}$ for $R$-LBMs. Data is gathered from lid-driven cavity simulations using $L=2^{13}$.
@@
Table 8: Overview of the mesh configurations for Aeolian noise simulations with the D2Q9-RR-LBM. Tree depths range from $14$ to $15$, with corresponding finest grid resolutions of $\Delta x_{f}\in\{D/40,D/80\}$, and total number of time iterations of $n_{\mathrm{ite}}\in\{$283’840, 567’616}\}.
@@
The consistent findings across all finest grid resolutions ($\Delta x_{f}\in\{D/20,D/40,D/80\}$) underline the importance of tailoring the grid to the specific simulation objectives. Whether the focus is on memory footprint, near-field aerodynamics, or far-field acoustics, adopting best practices to balance accuracy, performance, and simulation time remains crucial. The performance model developed here and the insights from the case studies offer a robust methodology for optimizing grid generation in a semi-automated manner, even for more complex scenarios requiring large domains or localized refinement patches.